OpenStudy (anonymous):

A wheel chair ramp rises 1.5ft every 8ft of horizontal distance it covers. Find the rate of change of the ramp.

OpenStudy (lgbasallote):

is this related rates? i mean calculus?

OpenStudy (anonymous):

\[\tan \theta=\frac{3}{16}=\frac{dy}{dx}\]

OpenStudy (anonymous):

@Igbasallote I don't know o.0 my book just said so.

OpenStudy (anonymous):

Read Your Text book That Would give you the most appropriate answer!

OpenStudy (anonymous):

I only have the txt book, I don't have the answer book, the answer book is with my teacher.

OpenStudy (anonymous):

o'well thanks for trying guyzz.